Show the frequencies of the notes in the different clarinet registers that are fingered nearly the same are in the ratio of 1 : 3 : 5.
Answer #1

let L be the length and v be the speed of sound in air

λ1 be the wavelength of first harmonic

λ2 be the wave length of third harmonic

λ3 be the wave length of fifth harmonic

now, λ1 = 4 L

hence, frequency of first harmonic f1 = v / λ1 = v / 4L

now, λ2 = 4L/3

frequency of third harmonic f2 = v / λ2 = v / (4L/3) = 3v / 4L

now, λ3 = 4L/5

frequency of fifth harmonic f3 = v / λ3 = v / (4L/5) = 5v / 4L

hence, f1 : f2 : f3 = v / 4L : 3v / 4L : 5v / 4L = 1 : 3 : 5